2. Assume that a researcher randomly selects 14 newborn babies and counts the number of (1 point) girls, x. The probabilities corresponding to the 14 possible values of x are summarized irn the given table. Answer the questions using the table. Probabilities of Girls xígirls) P) k(girls) Po) k(girls) Po) 0 0+ 5 0.122 10 0.061 1 0.001 6 0.183 11 0.022 2 0.006 7 0.209 12 0.006 3 0.022 80.18313 0.001 4 0.0619 0.122 14 0+ Find the combined probability of selecting 5, 6. 7. 8, or 9 girls. o0.122 0.610 00.819 0.989
Answer #1


Solution :

=> Answer :- 0.809

From given table , x(girls) as 5 , 6 , 7 , 8 , 9

P(x) as 0.122, 0.183, 0.209, 0.183, 0.122

=> The required combined probability is = P(x = 5) + P(x = 6) + P(x = 7) + P(x = 8) + P(x = 9)

= 0.122 + 0.183 + 0.209 + 0.183 + 0.122

= 0.809